OpenStudy (rcsd1):

Find the quotient. Simplify your answer. (n+8/n) divided by (7/n)

satellite73 (satellite73):

\[\frac{n+8}{n}\div \frac{7}{n}=\frac{n+8}{n}\times \frac{n}{7}\] jjust like with numbers

OpenStudy (triciaal):

when you divide by a fraction keep the dividend change the division to multiplication and turn the divisor upside down (invert/ flip)

OpenStudy (triciaal):